A 1200 kg car is traveling at 30 m/s. What is the kinetic energy of the car?
1
Expert's answer
2012-11-19T09:14:05-0500
KE = 1/2 * mV^2
KE = 1/2 * 1200 * 30^2
KE = 1/2 * 1200 * 900
KE = 540000